# Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)

### v2

@@ 4 Methodology
The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors $\eta_{0},\eta_{1},...,\eta_{m}$. This decomposition facilitates the derivation of the action-value function as a sum of delta estimators, where each estimator indicates the distinction between action-value functions at various discount factors. Concentrating on smaller, more manageable time scales facilitates the regulation of learning variance and bias, which constitutes the primary benefit of this approach.
@@ 4.2 Single-Step TD ( SARSA($\Delta$) )
This section initiates with an overview of the delta estimators ($D_{m}$) pertaining to the action-value functions $Q$ employed in SARSA. SARSA is extended to TD($\Delta$), which is referred to as SARSA($\Delta$). SARSA($\Delta$) seeks to derive the action-value function $Q(s,a)$, which indicates the anticipated return of following policy $\pi$ from state $s$, executing action $a$, and then persistently adhering to $\pi$ thereafter. The objective is to enhance the total reward by optimizing the action-value function over various time-scales. The fundamental concept in SARSA($\Delta$) is the segmentation of the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, where each component is linked to a distinct discount factor $\eta_{m}$. The relationship between these delta components (i.e., delta function) is expressed in the following manner:
@@
The update rule in standard SARSA is determined by the state-action pair in accordance with the policy. The action-value function is updated through the TD error, with the action in the subsequent state selected in accordance with the policy:
@@
In this context, $r_{n}$ signifies the reward acquired from performing action $a_{n}$ in state $s_{n}$, whereas $(s_{n+1},a_{n+1})$ denotes the ensuing state-action pair. The SARSA update rule is implemented independently to each delta component $D_{n}(s,a)$. The modifications for each component utilize the same architecture as the conventional SARSA approach, but it is tailored to include the delta function at each distinct time-scale. The update in Eq.11 for single-step TD SARSA ($\Delta$) can be articulated with numerous time-scales by employing the principle of splitting the update into several discount factors[8] as demonstrated below:
@@
where $Q_{\eta_{m}}(s_{n},a_{n})$ represents the action-value function with a discount factor $\eta_{m}$. $D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for action-value associated with the subsequent state-action pair. The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V. The single-step Bellman Eq. for the Q-value function in standard SARSA is as follows:
@@
This denotes the expected value of the reward acquired at time n, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{n+1},a_{n+1})$. Currently, rather than utilizing a singular discount factor $\eta$, we are employing a sequence of discount factors $\eta_{0},\eta_{1},.....,\eta_{m}$. The delta function $D_{m}(s_{n},a_{n})$ is defined as the difference between action-value functions associated with successive discount factors from Eq. 9:
@@
To expand the delta function $D_{z}$ utilizing its definition, we replace the Bellman Eq. for both $Q_{\eta_{m}}$ and $Q_{\eta_{m-1}}$, thereby expressing $D_{m}$ in relation to these action-value functions. Commence by composing:
@@
Decomposing the terms allows for a more detailed analysis of the expression. Utilizing the recursive relationship from Eq. 14
@@
This Eq. demonstrates that the update for $D_{m}$ integrates the disparity in discount factors $\eta_{m}-\eta_{m-1}$ with respect to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the subsequent step $D_{m}(s_{n+1},a_{n+1})$. This denotes a Bellman Eq. for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$. Consequently, it can be utilized to define the expected TD update for $D_{m}$. In this expression, $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ can be expressed as the summation of $D_{m}(s_{n+1},a_{n+1})$ for $m\leq M-1$, indicating that the Bellman Eq. for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale. Thus, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping. This process necessitates the assumption of a sequence of discount factors, denoted as $\eta_{m}$, which includes both the minimum and maximum values, $\eta_{0}$ and $\eta_{m}$[8].
@@ 4.3 Multi-Step TD ( SARSA($\Delta$) )
Numerous studies indicate that multi-step TD methods typically exhibit greater efficiency compared to single-step TD methods [29]. In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward. This approach takes into account the differences between consecutive discount factors, $\eta_{m}$ and $\eta_{m-1}$, while utilizing the value function at state $s_{n+k_{m}}$ for bootstrapping. In SARSA($\Delta$), the update Eqs. are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators. In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $D_{m}$.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (17)
In this case, $r_{m+x}$ represents the reward obtained at time step $n+x$. $a_{n+k_{m}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, where $k_{m}$ is the number of steps linked to the discount factor $\eta_{m}$. This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions. In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step. For example, utilizing a singular discount factor $\eta$, the multi-step TD update equation is expressed as:
@@
We now implement the multi-step variant of the Bellman Eq. for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$. The multi-step Bellman Eq. for $Q_{\eta_{m}}(s_{n},a_{n})$ is expressed as follows:
$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
@@
$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
We will now utilize the delta estimator formulas to subtract these two Eqs.
$D_{m}(s_{n},a_{n})=\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]-\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
We will now broaden and simplify the expressions. The immediate reward terms $r_{n+x}$ are present in both Eqs. but are adjusted by distinct discount factors, $\eta_{m}$ and $\eta_{m-1}$. This enables us to express the disparity in rewards as:
@@
Subsequently, for the bootstrapping terms, we utilize the recursive relationship as outlined:
@@
As a result, we get:
$\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
We now amalgamate the terms for both the rewards and bootstrapping to produce the final expression:
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (19)
This Eq. extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL. As a result, each $D_{m}$ obtains a share of the rewards from the environment up to time-step $k_{m}-1$. Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale. A variation of this algorithm, utilizing k-step bootstrapping as described in [29], is presented in Algorithm 1. Despite Algorithm 1 exhibiting quadratic complexity in relation to M, it can be executed with linear complexity for substantial M by preserving $\hat{Q}$ action-values at each time-scale $\eta_{m}$.
@@ 4.4 SARSA TD($\lambda,\Delta$)
Eq.20 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{n}^{\eta,\lambda}$ is defined as:
@@
Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function. Similarly, Eq.22 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
@@
where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})-\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
@@ 4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [28] seeks to compute the advantage function by aggregating multi-step TD errors. To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, we compute advantage estimates $A^{\Delta}(s_{n},a_{n})$ for each time-scale. Each advantage estimate $A^{\Delta}(s_{n},a_{n})$ employs a multi-step TD error pertinent to the delta estimator $D_{n}$, expressed as follows:
@@
We utilize $\eta_{m}$ as the discount factor and the sum of all D estimators as a surrogate for $Q_{\eta m}$. This objective is applicable to PPO by implementing the policy update from Eq.7 and replacing A with $A^{\Delta}$. Additionally, to train each $D_{m}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.22. For more information, see Algorithm 2.
@@
                  $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$triangleright Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq.22.
@@
              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq.23.
              $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$triangleright Update $\theta^{m}$ with TD (Eq.3) using $G^{m,\lambda_{m}}\forall m$.
              $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$triangleright from Eq.7
              $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$triangleright Update the policy parameters $\nu$ with PPO (Eq.7) for SARSA using $A^{\Delta}$

### v4

@@ 4 Methodology
The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors, $\eta_{0},\eta_{1},…\eta_{m}$. This breakdown makes it easier to find the action-value function as a sum of delta estimators, where each estimator shows how action-value functions differ at different discount factors. The main advantage of this method is that it makes it easier to control learning variance and bias by focusing on smaller, more controllable time scales.
@@ 4.2 Single-Step TD ( SARSA($\Delta$) )
This section initiates with an overview of the delta estimators ($D_{m}$) pertaining to the action-value functions $Q$ employed in SARSA. SARSA is expanded to TD($\Delta$), which is called SARSA($\Delta$). SARSA($\Delta$) aims to determine the action-value function $Q(s,a)$, representing the expected return of adhering to policy $\pi$ from state $s$, executing action $a$, and subsequently continuing to follow $\pi$. The goal is to optimize the total reward by improving the action-value function over various time scales. SARSA($\Delta$) fundamentally decomposes the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, each associated with a distinct discount factor $\eta_{m}$. The relationship between these delta components (i.e., delta function) is expressed in the following manner:
@@
The policy dictates how the state-action pair in conventional SARSA determines the update rule. The TD error adjusts the action-value function, and the policy determines the action in the following state:
@@
In this context, $r_{n}$ signifies the reward acquired from performing action $a_{n}$ in state $s_{n}$, whereas $(s_{n+1},a_{n+1})$ denotes the ensuing state-action pair. All of the delta components $D_{n}(s,a)$ are updated separately using the SARSA rule. The modifications to each component use the same structure as the standard SARSA method, but they are changed to incorporate the delta function at each time scale. The update in Eq. 11 for single-step TD SARSA ($\Delta$) can be expressed with numerous time-scales by utilizing the principle of splitting the update into several discount factors [1] as demonstrated below:
@@
In this context, $Q_{\eta_{m}}(s_{n},a_{n})$ denotes the action-value function with a discount factor ($\eta_{m}$). $D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for the action-value associated with the subsequent state-action pair. The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V. The single-step Bellman Eq. for the Q-value function in standard SARSA is as follows:
@@
This denotes the expected value of the reward acquired at time n, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{n+1},a_{n+1})$. Currently, instead of utilizing a single discount factor $\eta$, we are using a series of discount factors $\eta_{0},\eta_{1},...,\eta_{m}$. The delta function $D_{m}(s_{n},a_{n})$ is defined as the difference between action-value functions associated with successive discount factors from Eq. 9:
@@
To expand the delta function $D_{m}$ utilizing its definition, we replace the Bellman Eq. for both $Q_{\eta_{m}}$ and $Q_{\eta_{m-1}}$, thereby expressing $D_{m}$ in relation to these action-value functions. Commence by composing:
@@
Breaking down the terms facilitates a more thorough examination of the expression. Utilizing the recursive relationship from Eq. 14
@@
Eq. 4.2 shows that the difference in discount factors $\eta_{m}-\eta_{m-1}$ is integrated into the update for $D_{m}$ with regard to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the next step $D_{m}(s_{n+1},a_{n+1})$. The Eq. 4.2 denotes a Bellman Eq. 13 for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$. Consequently, it can be utilized to define the expected TD update for $D_{m}$. In this expression, $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ can be expressed as the summation of $D_{m}(s_{n+1},a_{n+1})$ for $m\leq M-1$, indicating that the Bellman Eq. 13 for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale. Consequently, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained concurrently through a TD update, employing the prior values of each estimator for bootstrapping. This process necessitates the assumption of a sequence of discount factors, denoted as $\eta_{m}$, which includes both the minimum and maximum values, $\eta_{0}$ and $\eta_{m}$ [1].
@@ 4.3 Multi-Step TD ( SARSA($\Delta$) )
Numerous studies indicate that multi-step TD methods typically exhibit greater efficiency compared to single-step TD methods [12]. In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward. This approach takes into account the differences between consecutive discount factors, $\eta_{m}$ and $\eta_{m-1}$, while utilizing the value function at state $s_{n+k_{m}}$ for bootstrapping. In SARSA($\Delta$), the update Eqs. are adjusted to combine rewards over a sequence of transitions before bootstrapping from either the Q-values or delta estimators. In multi-step SARSA($\Delta$), the agent accumulates rewards over K steps, resulting in a necessary adjustment to the update rule for each $D_{m}$.
$D_{m}(s_{n},a_{n})$ $=\mathbb{E}\biggl{[}\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$ (16)
$+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}$
In this case, $r_{n+x}$ represents the reward obtained at time step $n+x$. The term $a_{n+k_{m}}$ and the action are selected in a greedy manner as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, with $k_{m}$ representing the number of steps associated with the discount factor $\eta_{m}$. This approach extends Q-learning to multi-step temporal difference learning by gathering rewards over multiple stages and employing bootstrapping from both the present and previous time-scale action-value functions. In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step. For instance, employing a single discount factor $\eta$, the multi-step TD update Eq. can be articulated as:
@@
The multi-step variant of the Bellman Eq. is now implemented for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$. The multi-step Bellman Eq. for $Q_{\eta_{m}}(s_{n},a_{n})$ is expressed as follows:
$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}$
@@
$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}$
The formulas for the delta estimator are utilized to perform the subtraction of the two Eqs.
$\begin{split}D_{m}(s_{n},a_{n})&=\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}\\ &-\biggl{[}\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}\end{split}$
The expressions are broadened and simplified. The immediate reward terms $r_{n+x}$ are present in both Eqs. but are adjusted by distinct discount factors, $\eta_{m}$ and $\eta_{m-1}$. This enables us to express the disparity in rewards as:
@@
Subsequently, for the bootstrapping terms, the recursive relationship is utilized as outlined:
@@
As a result, the following expression is obtained:
$\begin{split}\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})&=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\\ &+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\end{split}$
The definitions for both the rewards and bootstrapping are now integrated to yield the final expression:
$\begin{split}D_{m}(s_{n},a_{n})&=\mathbb{E}\biggl{[}\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\\ &+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\biggr{]}\end{split}$ (18)
Eq. 18 extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL. As a result, each $D_{m}$ obtains a share of the rewards from the environment up to time-step $k_{m}-1$. Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time scale. A variation of this algorithm, utilizing k-step bootstrapping as described in [12], is presented in Algorithm 1. Despite Algorithm 1 exhibiting quadratic complexity in relation to M, it can be executed with linear complexity for substantial M by preserving $\hat{Q}$ action-values at each time-scale $\eta_{m}$.
@@ 4.4 SARSA TD($\lambda,\Delta$)
Eq. 19 presents the $\lambda$-return [34, 2], which integrates rewards across several steps to establish a target for TD($\lambda$) updates. The $\lambda$-return $G_{n}^{\eta,\lambda}$ is formally defined as follows:
@@
P represents the transition matrix for state-action pairs according to the policy $\pi$, while Q denotes the action-value function. Similarly, Eq. 21 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
@@
where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})\\ -\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
@@ 4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [39] seeks to compute the advantage function by aggregating multi-step TD errors. To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, advantage estimates $A^{\Delta}(s_{n},a_{n})$ are computed for each time scale. Each advantage estimate $A^{\Delta}(s_{n},a_{n})$ employs a multi-step TD error pertinent to the delta estimator $D_{n}$, expressed as follows:
@@
The discount factor $\eta_{m}$ is utilized, and the sum of all D estimators as a surrogate for $Q_{\eta m}$. This objective is applicable to PPO by implementing the policy update from Eq. 7 and replacing A with $A^{\Delta}$. Additionally, to train each $D_{m}$, a truncated form of their corresponding $\lambda$-return are used, as outlined in Eq. 21. For more information, see Algorithm 2.
@@
      $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$ $\triangleright$ Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq. 21.
@@
     $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$ $\triangleright$ Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq. 22.
     $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$ $\triangleright$ Update $\theta^{m}$ with TD (Eq. 3) using $G^{m,\lambda_{m}}\forall m$.
     $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$ $\triangleright$ from Eq. 7
     $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$ $\triangleright$ Update the policy parameters $\nu$ with PPO (Eq. 7) for SARSA using $A^{\Delta}$